# arXiv:2602.19846v2[math.AP] 02 Mar 2026

### v2

@@ 1.4. Organization of this paper
This paper is organized as follows. In Section 2, we summarize the notations used in this paper and elementary facts on Besov and Chemin–Lerner spaces. In Section 3, we provide key Besov estimates for oscillating functions, which play a key role in constructing the leading part of the non-unique stationary solutions. In Section 4, we construct the principal part of the non-trivial steady state and calculate Besov norms of it and the perturbative force. In Section 5, we make use of special flow constructed in Section 4 to complete the proof of Theorem 1.2.
@@ 3. Highly oscillating functions in Besov spaces
$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,k,\ell},$ (3.31)
@@
$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}\int_{\mathbb{R}^{n}}(f_{m}(x-y)-f_{m}(x))\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.60)
@@
${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{m}\right\|}_{L^{p}}-C\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.66)
@@ 4.1. Notations and key lemmas
    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \sqrt{2}\left(\mathbbm{e}_{i}\mp\mathbbm{e}_{j}\right),&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j).\end{cases}$ (4.8)
@@
    We note that, throughout this section, $\lambda$ may be replaced by a larger dyadic number whenever necessary, without further comment. We emphasize that the required lower bound on $\lambda$ is chosen independently of $j$. Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in their notations so as to avoid unnecessarily complicating them.
@@
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{4}{\left\|\nabla^{m+2}\rho_{k}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.24)
@@
${\left\|\Psi_{j,k}\right\|}_{L^{p}}\leqslant\sum_{\ell\in\Lambda_{j}}\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\leqslant\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{4\mu}_{j-1}\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{-4\mu}_{j-1}$ (4.26)
@@
${\left\|\mathcal{D}\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{6}\leqslant r_{0}\lambda_{j}^{8}$ (4.27)
@@
$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\frac{\lambda_{j-1}^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}.$ (4.73)
@@
$V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}={}2\lambda_{j-1}^{8}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}\otimes\Phi_{j,k_{2},\ell_{2}}$ (4.99)
@@
$\qquad+2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell}(x)\Phi_{j,k_{2},\ell}(x)$ (4.102)
$\quad={}\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j},$ (4.103)
@@
$F_{1,1;j}:=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right]^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.104)
$F_{1,2;j}:=2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x).$ (4.105)
@@
$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}\right)\right)$ (4.107)
$=\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,1;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,2;j}$ (4.108)
$=:F_{1,1}+F_{1,2}.$ (4.109)
@@
${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.116)
@@
Here, one may recall (3.13) for the definition of $\vartheta(\cdot,\cdot)$. Hence, we obtain
${\left\|F_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j=1}^{\infty}\left({\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\right)$ (4.123)
@@
$\leqslant{}$ $C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu}.$ (4.138)
@@
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{4-2\mu}$ (4.140)
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{8-(6-\frac{n}{r})\mu}$ (4.141)
$\leqslant C\lambda^{8+n-(6-\frac{n}{r})\mu}\leqslant C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.142)
@@
$\leqslant C\left(\lambda^{\frac{9}{2}-\frac{5}{2}\mu}+\lambda^{n+6-2\mu}\right)\lambda^{n+6-2\mu}$ (4.145)
@@
$\begin{cases}-\Delta W_{\lambda}+\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W_{\lambda}+W_{\lambda}\otimes V_{\lambda}+W_{\lambda}\otimes W_{\lambda}\right)=-F,\quad&x\in\mathbb{R}^{n},\\ \operatorname{div}W_{\lambda}=0,\quad&x\in\mathbb{R}^{n},\end{cases}$ (5.1)
@@
${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant 2C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
@@
$\begin{cases}\partial_{t}w_{\lambda}-\Delta w_{\lambda}+\mathbb{P}\operatorname{div}(U_{\lambda}\otimes w_{\lambda}+U_{\lambda}\otimes w_{\lambda}+w_{\lambda}\otimes w_{\lambda})=0,\\ \operatorname{div}w_{\lambda}=0,\\ w_{\lambda}(0,x)=w_{0}:=u_{0}-U_{\lambda}.\end{cases}$ (5.11)
@@
for all $w,w_{1},w_{2}\in\widetilde{C}([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some positive constant $C_{2}=C_{2}(n,p,r,\theta)$. We define the solution space $\mathscr{Y}$ by
@@
Then, we see that $\mathcal{Y}$ is a contraction mapping on $\mathscr{Y}$, by choosing $\lambda$ and ${\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}$ sufficiently large and small, respectively. Hence, the Banach fixed point theorem implies that there exists a unique mild solution $w_{\lambda}\in\mathscr{Y}$ to the perturbed system (5.11). We then obtain the non-stationary solution to (1.1) by $u_{\lambda}:=U_{\lambda}+w_{\lambda}$.

### v3

@@ 1.4. Organization of this paper
This paper is organized as follows. In Section 2, we summarize the notations used in this paper and elementary facts on Besov and Chemin–Lerner spaces. In Section 3, we provide key Besov estimates for oscillating functions, which play a key role in constructing the leading part of the non-unique stationary solutions. In Section 4, we construct the principal part of the non-trivial steady state and calculate its Besov norms and the perturbative force. In Section 5, we make use of the special flow constructed in Section 4 to complete the proof of Theorem 1.2.
@@ 3. Highly oscillating functions in Besov spaces
$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{k}\Delta_{j}g_{\lambda,k,\ell},$ (3.31)
@@
$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}\int_{\mathbb{R}^{n}}(f_{m}(x-y)-f_{m}(x))\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.60)
@@
${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{m}\right\|}_{L^{p}}-C\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{m}\right\|}_{L^{p}}.$ (3.66)
@@ 4.1. Notations and key lemmas
    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \dfrac{\mathbbm{e}_{i}\mp\mathbbm{e}_{j}}{\sqrt{2}},&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j).\end{cases}$ (4.8)
@@
    Throughout this section, $\lambda$ may be replaced by a larger dyadic number whenever necessary, without further comment. We emphasize that the required lower bound on $\lambda$ is chosen independently of $j$. Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in their notations so as to avoid unnecessarily complicating them. We also note that the generic constants are always independent of $\lambda$ and $j$.
@@
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{4}{\left\|\nabla^{m+2}\rho_{j}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.24)
@@
${\left\|\Psi_{j,k}\right\|}_{L^{p}}\leqslant\sum_{\ell\in\Lambda_{j}}\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\leqslant\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{4\mu}_{j-1}\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{n-4\mu}_{j-1}$ (4.26)
@@
${\left\|\mathcal{D}\mathcal{V}_{j}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}_{j}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{6}\leqslant r_{0}\lambda_{j}^{8}$ (4.27)
@@
$\leqslant\begin{cases}C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\left(\mu\log_{\lambda}\lambda_{j-1}\right)^{\frac{1}{q}-\frac{1}{2}}&(p=n,\ q>2),\\[5.0pt] C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\lambda_{j-1}^{(\frac{n}{p}-1)\mu}\left(\mu\log_{\lambda}\lambda_{j-1}\right)^{-\frac{1}{2}}&(p>n)\end{cases}$ (4.73)
$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}.$ (4.74)
@@
$V^{({\rm p})}_{j}(x)\otimes V^{({\rm p})}_{j}(x)={}2\lambda_{j-1}^{8}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x)$ (4.100)
@@
$\qquad+2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell}(x)\otimes\Phi_{j,k_{2},\ell}(x)$ (4.103)
$\qquad+2\lambda_{j-1}^{8}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell_{1},\ell_{2}\in\Lambda_{j}\\ \ell_{1}\neq\ell_{2}\end{subarray}}\Phi_{j,k,\ell_{1}}(x)\otimes\Phi_{j,k,\ell_{2}}(x)$ (4.104)
$\quad={}\lambda_{j-1}^{8}\psi_{j}(x)^{2}{\rm Id}-\mathcal{D}V_{j-1}(x)+F_{1,1;j}(x)+F_{1,2;j}(x)+F_{1,3;j}(x),$ (4.105)
@@
$F_{1,1;j}(x):=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right](x)^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.106)
$F_{1,2;j}(x):=2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x),$ (4.107)
$F_{1,3;j}(x):=2\lambda_{j-1}^{8}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell_{1},\ell_{2}\in\Lambda_{j}\\ \ell_{1}\neq\ell_{2}\end{subarray}}\Phi_{j,k,\ell_{1}}(x)\otimes\Phi_{j,k,\ell_{2}}(x).$ (4.108)
@@
$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}+F_{1,3;j}\right)\right)$ (4.110)
$=\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,1;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,2;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,3;j}$ (4.111)
$=:F_{1,1}+F_{1,2}+F_{1,3}.$ (4.112)
@@
${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}=\ell_{1}}^{\infty}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.119)
@@
Here, one may recall (3.13) for the definition of $\vartheta(\cdot,\cdot)$. We may treat $F_{1,3;j}$ similarly. Indeed, as there holds
$F_{1,3;j}(x)=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{\pm}\sum_{\ell_{1}\neq\ell_{2}}\sum_{k\in\mathcal{K}}$ $\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j,k}[V_{j-1}](x)^{2}$ (4.126)
$\psi_{j}(x)^{2}\cos\left((\lambda^{\ell_{1}}\pm\lambda^{\ell_{2}})a_{k}\cdot x\right)k\otimes k,$ (4.127)
we have
${\left\|F_{1,3;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}=\ell_{1}+1}^{\infty}$ (4.128)
$\left(\frac{\lambda^{(\frac{n}{r}-2)\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|(1\pm\lambda^{\ell_{1}-\ell_{2}})a_{k}|^{\frac{n}{r}-2}{\left\|\Gamma_{j,k}[V_{j-1}]^{2}\psi_{j}^{2}\right\|}_{L^{r}}\right$. (4.129)
$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|(1\pm\lambda^{\ell_{1}-\ell_{2}})a_{k}|^{-2}{\left\|\nabla^{n}(\Gamma_{j,k}[V_{j-1}]^{2}\psi_{j}^{2})\right\|}_{L^{r}\cap L^{1}}\right)$ (4.130)
$\leqslant{}$ $C\lambda_{j-1}^{4n+8-4\mu}.$ (4.132)
Hence, we obtain
${\left\|F_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j=1}^{\infty}\left({\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,3;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\right)$ (4.133)
@@
$\leqslant{}$ $C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu+2n}.$ (4.148)
@@
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{4-2\mu+2n}$ (4.150)
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{8+2n-(6-\frac{n}{r})\mu}$ (4.151)
$\leqslant C\lambda^{8+3n-(6-\frac{n}{r})\mu}\leqslant C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.152)
@@
$\leqslant C\left(\lambda^{5-2\mu}+\lambda^{n+6-2\mu}\right)\lambda^{n+6-2\mu}$ (4.155)
@@
$\begin{cases}-\Delta W_{\lambda}+\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W_{\lambda}+W_{\lambda}\otimes V_{\lambda}+W_{\lambda}\otimes W_{\lambda}\right)=-F_{\lambda},\quad&x\in\mathbb{R}^{n},\\ \operatorname{div}W_{\lambda}=0,\quad&x\in\mathbb{R}^{n},\end{cases}$ (5.1)
@@
${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{3}{2}C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
@@
$\begin{cases}\partial_{t}w_{\lambda}-\Delta w_{\lambda}+\mathbb{P}\operatorname{div}(U_{\lambda}\otimes w_{\lambda}+w_{\lambda}\otimes U_{\lambda}+w_{\lambda}\otimes w_{\lambda})=0,\\ \operatorname{div}w_{\lambda}=0,\\ w_{\lambda}(0,x)=w_{0}:=u_{0}-U_{\lambda}.\end{cases}$ (5.11)
@@
for all $w,w_{1},w_{2}\in\widetilde{C}([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some positive constant $C_{2}=C_{2}(n,p,q,\theta)$. We define the solution space $\mathscr{Y}$ by
@@
Choosing $\lambda$ so large and ${\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}$ so small that
$2C_{2}\left(\lambda^{-\alpha(p)}+4C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\right)\leqslant 1,$ (5.20)
we have
${\left\|\mathcal{Y}[w]\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}\leqslant\frac{3}{2}C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}},$ (5.21)
${\left\|\mathcal{Y}[w_{1}]-\mathcal{Y}[w_{2}]\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}\leqslant\frac{1}{2}{\left\|w_{1}-w_{2}\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.22)
for all $w,w_{1},w_{2}\in\mathscr{Y}$, which implies that $\mathcal{Y}$ is a contraction mapping on $\mathscr{Y}$. Hence, the Banach fixed point theorem implies that there exists a unique mild solution $w_{\lambda}\in\mathscr{Y}$ to the perturbed system (5.11). We then obtain the non-stationary solution to (1.1) by $u_{\lambda}:=U_{\lambda}+w_{\lambda}$.